For what values of p does the vertex of the parabola y=x^2 +2px+13 lie at a distance of 5 units from the origin

Grade:11

2 Answers

Arun
5 years ago
Dear student
 
it is not clear that distance should be towarda x axis or y axis
it is clear that it is upward parabola
 
now
 
its b/2a equals to 5
 
2p/2 = 5
 
p = 5
Akash Ranjan
5 years ago
Find the vertex ( -p, 13-p^2 ) . 
Now using distance formula from origin , equate the distance as 5 .( Distance can be 5 in any direction  so it's totally wrong to assume it's along X or Y axis.  
 
p^2 +(13-p^2)^2 = 25 
Solve to get p = 3,-3,4,-4